OpenStudy (anonymous):

Evaluate the following limit: lim x-> -4 (2x^2 + 7x - 4)/ (x^2+7x+12) Answer = 7 Thanks. @Mathematics

OpenStudy (anonymous):

I tried simplifying and then plugging -4 but i keep getting a fraction :S

OpenStudy (anonymous):

\[\frac{2x^2 + 7x - 4}{x^2+7x+12}=\frac{(x+4)(2x-1)}{(x+3)(x+4)}=\frac{2x+1}{x+3}\] I hope you can do it now yourself

OpenStudy (agreene):

did you type this correctly? because the answer isnt 7... It's 9. Use l'Hopital's rule.

OpenStudy (anonymous):

\[\frac{2\cdot(-4)+1}{-4+3}=\frac{-8+1}{-1}=7\]

OpenStudy (anonymous):

Yeah I think I typed it out right, thanks a lot for your help Tomas & Agreene, I appreciate it.

OpenStudy (agreene):

Wolfram says it's 9... and did l'hopitals like I did O_o http://www.wolframalpha.com/input/?i=lim+x-%3E-4+%28%282x%5E2+%2B+7x+-+4%29%2F+%28x%5E2%2B7x%2B12%29%29

OpenStudy (anonymous):

Oh thats weird :S

OpenStudy (agreene):

I call shenanigans on this question!

OpenStudy (anonymous):

lol, i made mistake that's why i got 7... sorry, maybe it's correct version of this :DD

OpenStudy (anonymous):

Haha wait, I think you're right, it should be 2x-1 / x+3

OpenStudy (anonymous):

So then -8-1/-4+3 then = 9 :D

OpenStudy (anonymous):

I guess I have a retarded answer sheet

OpenStudy (agreene):

oh, i didn't spot that either! no shenanigans, just two methods to solve ^_^

OpenStudy (anonymous):

:D